# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework

### v2

@@ Appendix G Discussion
To jointly capture temporal dynamics and cross-channel dependencies under practical settings, we proposed a unified attention masking strategy that merges intra- and inter-channel attention into a single attention layer. This design avoids the need for separate attention modules, reducing the number of parameters and simplifying the overall model architecture. This parameter efficiency is particularly beneficial for datasets with limited data, as it reduces the risk of overfitting while maintaining the model’s ability to capture essential temporal and cross-channel patterns. However, this unified structure introduces scalability challenges when applied to high-dimensional datasets. As the number of channels increases, the attention layer must process a quadratic number of token interactions within a single computation step, leading to substantial computational overhead and slower training. A promising direction to mitigate this issue is to optimize the attention computation, for example, by decomposing the attention map or distributing the computation across multiple smaller attention modules.
Beyond scalability, another limitation lies in the objective function. Although our model demonstrates strong predictive performance across diverse practical settings, it still suffers from amplitude attenuation—especially over longer forecasting horizons—due to the use of a standard mean squared error loss. To address this, it is crucial to explore alternative loss functions that better preserve spectral characteristics and periodicity, rather than relying solely on point-wise accuracy. This issue also extends to evaluation. Conventional metrics such as MSE may fail to penalize overly smoothed forecasts, often favoring models that suppress meaningful high-frequency or transient patterns. As a result, structurally inaccurate predictions may still appear favorable under MSE, particularly in long-term forecasting scenarios dominated by coarse trends. Refining evaluation metrics to better capture structural fidelity is therefore essential. Designing both loss functions and evaluation protocols that align with the characteristics of asynchronous, partially observed time series remains an important and open research challenge.

### v4

@@ Appendix G Discussion
To jointly capture temporal dynamics and cross-channel dependencies under practical settings, we proposed a unified attention masking strategy that merges intra- and inter-channel attention into a single attention layer. This design avoids the need for separate attention modules, reducing the number of parameters and simplifying the overall model architecture. This parameter efficiency is particularly beneficial for datasets with limited data, as it reduces the risk of overfitting while maintaining the model’s ability to capture essential temporal and cross-channel patterns. However, this unified structure introduces scalability challenges when applied to high-dimensional datasets. As the number of channels increases, the attention layer must process a quadratic number of token interactions within a single computation step, leading to substantial computational overhead and slower training. A promising direction to mitigate this issue is to optimize the attention computation, for example, by explicit channel sparsification or grouped channel representations.
Beyond scalability, another limitation lies in the objective function. Although our model demonstrates strong predictive performance across diverse practical settings, it still suffers from amplitude attenuation, especially over longer forecasting horizons, due to the use of a standard mean squared error loss. To address this, it is crucial to explore alternative loss functions that better preserve spectral characteristics and periodicity, rather than relying solely on point-wise accuracy. This issue also extends to evaluation. Conventional metrics such as MSE may fail to penalize overly smoothed forecasts, often favoring models that suppress meaningful high-frequency or transient patterns. As a result, structurally inaccurate predictions may still appear favorable under MSE, particularly in long-term forecasting scenarios dominated by coarse trends. Refining evaluation metrics to better capture structural fidelity is therefore essential. Designing both loss functions and evaluation protocols that align with the characteristics of asynchronous, partially observed time series remains an important and open research challenge.